### v2

@@ 5. Some Preliminary Results
The first lemma states a rule for interchanging upper limits. A proof can be found in [4, Corollary 1.16].
@@
The second lemma provides sufficient conditions for the continuity of the Fenchel conjugate with respect to the norm topology in the dual space. This result follows directly from [28, Corollary 4.4.11].
@@
The next lemma concerns the approximation of a convex proper lower semicontinuous function by its Moreau envelope, together with its continuity properties. We refer to [76, Theorem 3.265] and [28, Theorem 4.4.17] for more details.
@@
In this subsection, we establish Theorem 4.1 in the case where $X$ is separable. Proof. Suppose that $X$ is a separable Banach space and denote $f:=\Gamma\text{-}\displaystyle\liminf_{n}f_{n}$. From Lemma 6.1, we already know that
@@
First, by [64, Theorem 3], we have
@@
where in the last equality we have used that $\left(\frac{\lambda}{2}\left\lVert\cdot\right\rVert^{2}\right)^{\ast}=\frac{1}{2\lambda}\left\lVert\cdot\right\rVert_{\ast}^{2}$. Next, due to the fact that ${\rm dom}\kern 1.19995pt\left(f_{n}+\frac{\lambda}{2}\left\lVert\cdot\right\rVert^{2}\right)\cap{\rm dom}\kern 1.19995pt\delta_{B_{k}}\neq\emptyset$ (recall that $a_{n}\in B_{k}$), we may apply [36, Proposition 4.1.16] to deduce
@@
Note that by [36, Proposition 4.1.20], we have
@@ 6.2. Separable reduction and rich families
Separable reducibility is a common tool in functional analysis, as it often leads to substantial simplifications and, in some cases, provides the only viable approach to proving certain results. A simple example of a separably reducible property is continuity of a mapping. For a list of some less trivial properties that are separable determined in analysis, we refer the reader to [44]. There are special families of separable subspaces of a given Banach space that allow the implementation of a separable reduction, that is, the possibility to “lift” a property from separable subspaces to the entire space, since they permit countable many different properties to be dealt with. These special families are called rich families, and they were introduced by J.M Borwein and W.B Moors in [26]. Rich families were deeply discussed in connection with a number of properties related to differentiability and subdifferentiability in [53] and [45].
@@
The main property and the power of rich families is shown in the following fundamental fact (see [26, Proposition 1.1]).
@@ 6.4. Preliminary Results on Separable Reduction Techniques
Indeed, let $V\times Y\in\mathcal{R}$, and let $x^{\ast}\in\overline{Y}^{w^{\ast}}$. By the Mackey–Arens theorem (see, e.g., [46, Lemma 19]), there exists a net $(x^{\ast}_{\alpha})\subset Y$ converging to $x^{\ast}$ with respect to the Mackey topology, that is, $x^{\ast}_{\alpha}\overset{\mu^{\ast}}{\to}x^{\ast}$. Moreover, from [36, Proposition 4.1.20], we have
@@
Despite the fact that the previous result will be primarily used in the proof of Theorem 4.1, it is worth noting that it also leads to interesting insights into the structure of WCG Banach spaces. In particular, it provides a simple isometry between separable subspaces and the weak-closures of separable subspaces in the dual. This result is in the same spirit as [37, Theorem 2.3], but formulated for WCG spaces rather than Asplund spaces. Formally, we obtain the following result.
@@
By the Brøndsted–Rockafellar theorem (see, e.g.,  [30] or [28, Theorem 4.3.2]), for every $n\in\mathbb{N}$ one can find $(x_{n},y_{n}^{\ast})\in\partial f_{n}$ such that
@@
In order to prove that $(b)\Longrightarrow(c)$ in Theorem 4.7, we will recall the following integration Lemma given in [32, Lemma 2.3]. For that purpose let us introduce the following notation. For a function $\varphi\colon X\to\mathbb{R}\cup\{-\infty,\infty\}$ (not necessarily convex), the Fenchel subdifferential of $\varphi$ at a point $x\in X$ is defined as
@@
The following technical lemma will be useful in the sequel. Since its proof is only a slight modification of [32, Lemma 2.4], we omit it here.

### v3

@@ 4. Limits of convex functions and subdifferentials
5. Application to convex integral functionals
In this section we present an application of Corollary 4.6 to convex integral functionals. Functionals of this type arise naturally in the calculus of variations, optimal control, and stochastic optimization, and were systematically studied by Rockafellar in [71]. Let $(T,\mathcal{F},\mu)$ be a $\sigma$-finite measure space. Let $1\leq p<\infty$ and consider $L$ to be the space $L_{n}^{p}(T)$ consisting of all $\mathbb{R}^{n}$-valued measurable functions $u$ on $T$ such that
$\int_{T}|u(t)|^{p}d\mu(t).$
With $|\cdot|$ denoting the euclidean norm in $\mathbb{R}^{n}$. We will consider functionals of the form
$I_{f}(u):=\int_{T}f(t,u(t))d\mu(t)\text{ for }u\in L,$
where $f$ is a function from $T\times\mathbb{R}^{n}$ to $\mathbb{R}\cup\{\infty\}$ such that $f$ is convex integrand, i.e, $f(t,x)$ is a convex function of $x\in\mathbb{R}^{n}$ for each $t\in T$, hence $I_{f}$ is convex. Moreover, we will consider $f$ to be a normal integrand, i.e, $f(t,\cdot)$ is proper and lower semicontinuous for every $t\in T$ and if further there exists a countable collection $U$ of measurable functions $u$ from $T$ to $\mathbb{R}^{n}$ having the following properties:
  1. (1)
    for each $u\in U$, $f(t,u(t))$ is measurable in $t$;
  2. (2)
    for each $t$, $U_{t}\cap\operatorname{dom}f_{t}$ is dense in $\operatorname{dom}f_{t}$, where
    $U_{t}=\{u(t):u\in U\}.$
Although these last assumptions might seem complicated, there are some remarkable examples of normal integrands mentioned in [71].
Example 5.1.
Suppose $f(t,x)=F(x)$ for all $t\in T$, where $F$ is a lower semicontinuous proper convex function on $\mathbb{R}^{n}$. Then $f$ is a normal convex integrand.
Example 5.2.
Suppose $f$ is a convex integrand such that $f(t,x)$ is measurable in $t$ for each fixed $x\in\mathbb{R}^{n}$, and such that, for every $t\in T$, $f(t,x)$ is lower semi-continuous in $x$ and has interior points in its effective domain $\{x:f(t,x)<+\infty\}$. Then $f$ is a normal convex integrand. In particular, if $f$ is a convex integrand having only finite values such that $f(\cdot,x)$ is measurable for each $x$, the $f$ is a normal convex integrand.
We recall from [71, Lemma 5] that if $f$ is a normal convex integrand then $f^{\ast}$ is also a normal convex integrand and $f(t,u(t))$ is a measurable function of $t$ for every measurable function $u$ from $T$ to $\mathbb{R}^{n}$.
Lastly, we recall the following theorem and corollary from [71, Theorem 2].
Theorem 5.3.
Let $f$ be a normal convex integrand such that $f(t,u(t))$ is $\mu$-integrable for at least one $u\in L$, and $f^{\ast}(t,u^{\ast}(t))$ is $\mu$-integrable for at least one $u^{\ast}\in L^{\ast}$. Then $I_{f}$ on $L$ and $I_{f^{\ast}}$ on $L^{\ast}$ are proper convex functions conjugate to each other.
Corollary 5.4.
Suppose that $T$ is of finite measure. Let $f$ be of the form $f(t,x)=F(x)$, where $F$ is a lower semicontinuous proper convex function on $\mathbb{R}^{n}$. Then $I_{f}$ on $L$ and $I_{f^{\ast}}$ on $L^{\ast}$ are conjugate to each other.
Since $\mu$ is $\sigma$-finite, the space $L^{p}_{n}(T)$ is WCG for every $p\in[1,\infty)$, regardless of separability (which can not be ensured), and this includes the non-reflexive case $p=1$. Consequently, Corollary 4.6 applies to the convex integral functionals considered in this section.
Corollary 5.5.
Let $f,f_{n}\colon T\times\mathbb{R}^{n}\to\mathbb{R}\cup\{\infty\}$, $n\in\mathbb{N}$, be normal convex integrands. Assume that there are $u,u_{n}\in L$, $n\in\mathbb{N}$, and $u^{\ast},u^{\ast}_{n}\in L^{\ast},n\in\mathbb{N}$, such that $f(t,u(t)),f_{n}(t,u_{n}(t)),f^{\ast}(t,u^{\ast}(t)),f^{\ast}_{n}(t,u^{\ast}_{n}(t))$ are $\mu$-integrable. Suppose that
$\forall(v^{\ast}_{n})_{n\in\mathbb{N}}\subset L^{\ast},\sup_{n\in\mathbb{N}}I_{f^{\ast}_{n}}(v^{\ast}_{n})<\infty\implies\sup_{n\in\mathbb{N}}\left\lVert v^{\ast}_{n}\right\rVert_{\ast}<\infty.$
Then, $I_{f}=\Gamma\text{-}\displaystyle\lim_{n}I_{f_{n}}$ if and only if $I_{f^{\ast}}={\Gamma(w^{\ast})\text{-}\displaystyle\lim_{n}}I_{f^{\ast}_{n}}.$
Corollary 5.6.
Let $F,F_{n}\colon\mathbb{R}^{n}\to\mathbb{R}\cup\{\infty\}$, $n\in\mathbb{N}$, be proper lower semicontinuous convex functions. Assume that $T$ is of finite measure. In addition, suppose that
$\forall(v^{\ast}_{n})_{n\in\mathbb{N}}\subset L^{\ast},\sup_{n\in\mathbb{N}}I_{F^{\ast}_{n}}(v^{\ast}_{n})<\infty\implies\sup_{n\in\mathbb{N}}\left\lVert v^{\ast}_{n}\right\rVert_{\ast}<\infty.$
Then, $I_{F}=\Gamma\text{-}\displaystyle\lim_{n}I_{F_{n}}$ if and only if $I_{F^{\ast}}={\Gamma(w^{\ast})\text{-}\displaystyle\lim_{n}}I_{F^{\ast}_{n}}.$
When the sequence of functionals is locally equibounded, we may apply Corollary 4.8.
Corollary 5.7.
Let $f,f_{n}\colon T\times\mathbb{R}^{n}\to\mathbb{R}\cup\{\infty\}$, $n\in\mathbb{N}$, be normal convex integrands. Assume that there are $u,u_{n}\in L$, $n\in\mathbb{N}$, and $u^{\ast},u^{\ast}_{n}\in L^{\ast},n\in\mathbb{N}$, such that $f(t,u(t)),f_{n}(t,u_{n}(t)),f^{\ast}(t,u^{\ast}(t)),f^{\ast}_{n}(t,u^{\ast}_{n}(t))$ are $\mu$-integrable. Suppose that the sequence of functionals $(I_{f_{n}})_{n\in\mathbb{N}}$ is locally equibounded, i.e, for every $u\in L$, there exists $r>0$ and $M\in\mathbb{R}$ such that
$\forall n\in\mathbb{N},\forall v\in\mathbb{B}_{r}(u),\;|I_{f_{n}}(v)|\leq M.$
Then, the following are equivalent
  1. (a)
    $I_{f}=\Gamma\text{-}\displaystyle\lim_{n}I_{f_{n}}$.
  2. (b)
    $I_{f^{\ast}}={\Gamma(w^{\ast})\text{-}\displaystyle\lim_{n}}I_{f^{\ast}_{n}}$.
  3. (c)
    $I_{f}$ is the pointwise limit of $I_{f_{n}}$.
Corollary 5.8.
Let $F,F_{n}\colon\mathbb{R}^{n}\to\mathbb{R}\cup\{\infty\}$, $n\in\mathbb{N}$, be proper lower semicontinuous convex functions. Assume that $T$ is of finite measure, In addition, suppose that the sequence of functionals $(I_{f_{n}})_{n\in\mathbb{N}}$ is locally equibounded, i.e, for every $u\in L$, there exists $r>0$ and $M\in\mathbb{R}$ such that
$\forall n\in\mathbb{N},\forall v\in\mathbb{B}_{r}(u),\;|I_{f_{n}}(v)|\leq M.$
Then, the following are equivalent
  1. (a)
    $I_{f}=\Gamma\text{-}\displaystyle\lim_{n}I_{f_{n}}$.
  2. (b)
    $I_{f^{\ast}}={\Gamma(w^{\ast})\text{-}\displaystyle\lim_{n}}I_{f^{\ast}_{n}}$.
  3. (c)
    $I_{f}$ is the pointwise limit of $I_{f_{n}}$.
In what remains of this section, we will assume that $p\in(1,\infty)$, hence, L is reflexive, therefore, besides the two previous corollaries, we can apply Corollary 4.9 to deduce the following corollaries.
Corollary 5.9.
Let $f,f_{n}\colon T\times\mathbb{R}^{n}\to\mathbb{R}\cup\{\infty\}$, $n\in\mathbb{N}$, be normal convex integrands. Assume that there are $u,u_{n}\in L$, $n\in\mathbb{N}$, and $u^{\ast},u^{\ast}_{n}\in L^{\ast},n\in\mathbb{N}$, such that $f(t,u(t)),f_{n}(t,u_{n}(t)),f^{\ast}(t,u^{\ast}(t)),f^{\ast}_{n}(t,u^{\ast}_{n}(t))$ are $\mu$-integrable. Suppose that
$\forall(v_{n})_{n\in\mathbb{N}}\subset L,\sup_{n\in\mathbb{N}}I_{f_{n}}(v_{n})<\infty\implies\sup_{n\in\mathbb{N}}\left\lVert v_{n}\right\rVert<\infty.$
Then, $I_{f}={\Gamma(w)\text{-}\displaystyle\lim_{n}}I_{f_{n}}$ if and only if $I_{f^{\ast}}=\Gamma\text{-}\displaystyle\lim_{n}I_{f^{\ast}_{n}}.$
Corollary 5.10.
Let $F,F_{n}\colon\mathbb{R}^{n}\to\mathbb{R}\cup\{\infty\}$, $n\in\mathbb{N}$, be proper lower semicontinuous convex functions. Assume that $T$ is of finite measure. In addition, suppose that
$\forall(v_{n})_{n\in\mathbb{N}}\subset L,\sup_{n\in\mathbb{N}}I_{F_{n}}(v_{n})<\infty\implies\sup_{n\in\mathbb{N}}\left\lVert v_{n}\right\rVert<\infty.$
Then, $I_{F}={\Gamma(w)\text{-}\displaystyle\lim_{n}}I_{F_{n}}$ if and only if $I_{F^{\ast}}=\Gamma\text{-}\displaystyle\lim_{n}I_{F^{\ast}_{n}}.$
Lastly, in the case where the conjugates are locally equibounded, we can apply Corollary 4.10 to obtain the following corollaries.
Corollary 5.11.
Let $f,f_{n}\colon T\times\mathbb{R}^{n}\to\mathbb{R}\cup\{\infty\}$, $n\in\mathbb{N}$, be normal convex integrands. Assume that there are $u,u_{n}\in L$, $n\in\mathbb{N}$, and $u^{\ast},u^{\ast}_{n}\in L^{\ast},n\in\mathbb{N}$, such that $f(t,u(t)),f_{n}(t,u_{n}(t)),f^{\ast}(t,u^{\ast}(t)),f^{\ast}_{n}(t,u^{\ast}_{n}(t))$ are $\mu$-integrable. Suppose that the sequence of functionals $(I_{f^{\ast}_{n}})_{n\in\mathbb{N}}$ is locally equibounded, i.e, for every $u^{\ast}\in L^{\ast}$, there exists $r>0$ and $M\in\mathbb{R}$ such that
$\forall n\in\mathbb{N},\forall v^{\ast}\in\mathbb{B}^{\ast}_{r}(u^{\ast}),\;|I_{f^{\ast}_{n}}(v^{\ast})|\leq M,$
where $\mathbb{B}^{\ast}_{r}(u^{\ast})$ is the open ball of center $u^{\ast}$ and radius $r$ with the dual norm in $L^{\ast}$. Then, the following are equivalent
  1. (a)
    $I_{f^{\ast}}=\Gamma\text{-}\displaystyle\lim_{n}I_{f^{\ast}_{n}}$.
  2. (b)
    $I_{f}={\Gamma(w)\text{-}\displaystyle\lim_{n}}I_{f_{n}}$.
  3. (c)
    $I_{f^{\ast}}$ is the pointwise limit of $I_{f^{\ast}_{n}}$.
Corollary 5.12.
Let $F,F_{n}\colon\mathbb{R}^{n}\to\mathbb{R}\cup\{\infty\}$, $n\in\mathbb{N}$, be proper lower semicontinuous convex functions. Assume that $T$ is of finite measure. In addition, Suppose that the sequence of functionals $(I_{F^{\ast}_{n}})_{n\in\mathbb{N}}$ is locally equibounded, i.e, for every $u^{\ast}\in L^{\ast}$, there exists $r>0$ and $M\in\mathbb{R}$ such that
$\forall n\in\mathbb{N},\forall v^{\ast}\in\mathbb{B}^{\ast}_{r}(u^{\ast}),\;|I_{F^{\ast}_{n}}(v^{\ast})|\leq M,$
where $\mathbb{B}^{\ast}_{r}(u^{\ast})$ is the open ball of center $u^{\ast}$ and radius $r$ with the dual norm in $L^{\ast}$. Then, the following are equivalent
  1. (a)
    $I_{F^{\ast}}=\Gamma\text{-}\displaystyle\lim_{n}I_{F^{\ast}_{n}}$.
  2. (b)
    $I_{F}={\Gamma(w)\text{-}\displaystyle\lim_{n}}I_{F_{n}}$.
  3. (c)
    $I_{F^{\ast}}$ is the pointwise limit of $I_{F^{\ast}_{n}}$.
These results provide a dual characterization of the $\Gamma$-convergence of convex integral functionals on $L^{p}$ spaces in terms of the $\Gamma$-convergence of their conjugates with the $w^{\ast}$-topology.
@@ 6. Some Preliminary Results
The first lemma states a rule for interchanging upper limits. A proof can be found in [11, Corollary 1.16].
@@
The second lemma provides sufficient conditions for the continuity of the Fenchel conjugate with respect to the norm topology in the dual space. This result follows directly from [29, Corollary 4.4.11].
@@
The next lemma concerns the approximation of a convex proper lower semicontinuous function by its Moreau envelope, together with its continuity properties. We refer to [75, Theorem 3.265] and [29, Theorem 4.4.17] for more details.
@@
In this subsection, we establish Theorem 4.1 in the case where $X$ is separable.
Proof. Suppose that $X$ is a separable Banach space and denote $f:=\Gamma\text{-}\displaystyle\liminf_{n}f_{n}$. From Lemma 7.1, we already know that
@@
First, by [70, Theorem 3], we have
@@
where in the last equality we have used that $\left(\frac{\lambda}{2}\left\lVert\cdot\right\rVert^{2}\right)^{\ast}=\frac{1}{2\lambda}\left\lVert\cdot\right\rVert_{\ast}^{2}$. Next, due to the fact that ${\rm dom}\kern 1.19995pt\left(f_{n}+\frac{\lambda}{2}\left\lVert\cdot\right\rVert^{2}\right)\cap{\rm dom}\kern 1.19995pt\delta_{B_{k}}\neq\emptyset$ (recall that $a_{n}\in B_{k}$), we may apply [37, Proposition 4.1.16] to deduce
@@
Note that by [37, Proposition 4.1.20], we have
@@ 7.2. Separable reduction and rich families
Separable reducibility is a common tool in functional analysis, as it often leads to substantial simplifications and, in some cases, provides the only viable approach to proving certain results. A simple example of a separably reducible property is continuity of a mapping. For a list of some less trivial properties that are separable determined in analysis, we refer the reader to [45]. There are special families of separable subspaces of a given Banach space that allow the implementation of a separable reduction, that is, the possibility to “lift” a property from separable subspaces to the entire space, since they allow one to deal with countable many different hypotheses. These special families are called rich families, and they were introduced by J.M Borwein and W.B Moors in [27]. Rich families were deeply discussed in connection with a number of properties related to differentiability and subdifferentiability in [54] and [46].
@@
The main property and the power of rich families is shown in the following fundamental fact (see [27, Proposition 1.1]).
@@ 7.4. Preliminary Results on Separable Reduction Techniques
Indeed, let $V\times Y\in\mathcal{R}$, and let $x^{\ast}\in\overline{Y}^{w^{\ast}}$. By the Mackey–Arens theorem (see, e.g., [47, Lemma 19]), there exists a net $(x^{\ast}_{\alpha})\subset Y$ converging to $x^{\ast}$ with respect to the Mackey topology, that is, $x^{\ast}_{\alpha}\overset{\mu^{\ast}}{\to}x^{\ast}$. Moreover, from [37, Proposition 4.1.20], we have
@@
Despite the fact that the previous result will be primarily used in the proof of Theorem 4.1, it is worth noting that it also leads to interesting insights into the structure of WCG Banach spaces. In particular, it provides a simple isometry between separable subspaces and the weak-closures of separable subspaces in the dual. This result is in the same spirit as [38, Theorem 2.3], but formulated for WCG spaces rather than Asplund spaces. Formally, we obtain the following result.
@@
By the Brøndsted–Rockafellar theorem (see, e.g.,  [31] or [29, Theorem 4.3.2]), for every $n\in\mathbb{N}$ one can find $(x_{n},y_{n}^{\ast})\in\partial f_{n}$ such that
@@
In order to prove that $(b)\Longrightarrow(c)$ in Theorem 4.7, we will recall the following integration Lemma given in [33, Lemma 2.3]. For that purpose let us introduce the following notation. For a function $\varphi\colon X\to\mathbb{R}\cup\{-\infty,\infty\}$ (not necessarily convex), the Fenchel subdifferential of $\varphi$ at a point $x\in X$ is defined as
@@
The following technical lemma will be useful in the sequel. Since its proof is only a slight modification of [33, Lemma 2.4], we omit it here.